OpenStudy (anonymous):

explain the expression bellow and how to get the form : sin a + sin b = 2 sin (a+b)/2 cos (a-b)/2

OpenStudy (anonymous):

a=a+b/2+a-b/2 b=a+b/2-a-b/2 so you reach to your solution

OpenStudy (anonymous):